OpenStudy (anonymous):

How many 3*3*3 cubes will a 3*15*15 box hold?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

need explanation?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

3 / 3= 1 15/3=5 15/3=5 1 * 5*5=25

OpenStudy (anonymous):

thank you